OpenStudy (trojanpoem):

(y')^2 - 2xy' + (x-3)^2 = 0 Find the values of x whiches gives real solution to the D.E I don't even know what the question is asking for..

OpenStudy (holsteremission):

You might be able to solve for \(y'\) by completing the square/quadratic formula. Then the solution will probably involve some root term that has no imaginary component for some range of \(x\). I think it's this range that you're looking for.

OpenStudy (trojanpoem):

Let's try

OpenStudy (trojanpoem):

(y' - x)^2 - x^2 + (x-3)^2 = 0 (y'-x)^2 = x^2 - (x-3)^2 (y'-x)^2 = x^2 - (x^2 + 9 - 6x) = x^2 - x^2 - 9 + 6x (y'-x)^2 = 6x - 9 y'-x = sqrt(6x - 9) 6x- 9 >= 0 6x >= 9 x >= 3/2

OpenStudy (trojanpoem):

Respect bro,

OpenStudy (holsteremission):

Looks good to me. Happy to help :)

OpenStudy (trojanpoem):

Thx thx ^^